Show that π‘ˆ(βˆ’π‘“, 𝑝) = βˆ’πΏ(𝑓, 𝑝) and 𝐿(βˆ’π‘“, 𝑝) = βˆ’π‘ˆ(𝑓, 𝑝)


1
Expert's answer
2021-09-30T04:53:36-0400

Suppose that f is integrable on [a; b]. For "\in > 0" , there exist partitions P1 and P2 such that

"L(f, P_1) > L(f) − \frac{\in}{2} \space 
and \space U(f; P2) < U(f) + \frac{\in}{2}\\
For \space P := P_1 ∪ P_2 \space we \space have\\
L(f) −\frac{\in}{2} 
< L(f, P_1) ≤ L(f; P ) ≤ U(f, P ) ≤ U(f, P_2) < U(f) + \frac{\in}{2} \\"

Since L(f) = U(f), it follows that "U(f, P ) - L(f,P )< \in"

Implies π‘ˆ(βˆ’π‘“, 𝑝) = βˆ’πΏ(𝑓, 𝑝) and 𝐿(βˆ’π‘“, 𝑝) = βˆ’π‘ˆ(𝑓, 𝑝)